# Nonradial oscillations of stratified neutron stars with solid crusts: Mode characterization and tidal resonances in coalescing binaries

### v1

@@ A.1.2 Perturbation equations for elastic solid
$W^{\prime}$ $=-\frac{\ell+1}{r}W+re^{\lambda/2}\left[\frac{e^{-\nu/2}}{\Gamma_{1}p}X-\frac{2(n+1)}{r^{2}}V+\frac{1}{2}H_{2}+K\right]\,$, (73)
@@
$T_{2}^{\prime}$ $=-\frac{1}{2}re^{\lambda}(\epsilon+p)H_{0}+re^{\lambda-\nu/2}\left(X+\frac{1}{2r^{2}}e^{\nu/2}T_{1}\right)+\left[\frac{1}{2}(\lambda^{\prime}-\nu^{\prime})-\frac{\ell+1}{r}\right]T_{2}+\left[\frac{2ne^{\lambda}\mu}{r}-e^{\lambda-\nu}r\omega^{2}(\epsilon+p)\right]V+e^{\lambda/2}p^{\prime}W\,.$ (75)
@@
$\left[\frac{re^{-\lambda}}{2}\left(r\nu^{\prime}-2\right)+(n+1)r\right]H_{0}$ $=r^{2}e^{-\lambda}\left[\omega^{2}re^{-\nu}-\frac{n_{\ell}+1}{2}\nu^{\prime}\right]H_{1}$
@@
$\frac{1}{3}e^{-\nu/2}\mu r^{2}X+\frac{1}{4}\Gamma_{1}pT_{1}$ $=\frac{\mu\Gamma_{1}p}{2}\left[2e^{-\lambda/2}W-r^{2}K+2(n+1)V\right].$ (78)
@@ A.2 Joint and boundary conditions
The solutions must remain regular at the center of the star, necessitating a Taylor expansion of the perturbation variables near the origin to approximate the ODE solutions. It has been shown in Ref. [182] that not all the quantities are independent. Once $K(r=0)$ and $W(r=0)$ are specified, all other coefficients in the Taylor expansion can be uniquely determined (see details in Ref. [185, pp. 27–30, 59–60]). The surface of the star is a matter-vacuum interface. If a two-layer NS with a solid surface is considered, the boundary conditions at the surface are

### v2

@@ A.1.2 Perturbation equations for elastic solid
$W^{\prime}$ $=-\frac{\ell+1}{r}W+re^{\lambda/2}\left[\frac{e^{-\nu/2}}{\Gamma_{1}p}X-\frac{2(n_{\ell}+1)}{r^{2}}V+\frac{1}{2}H_{2}+K\right]\,$, (74)
@@
$T_{2}^{\prime}$ $=-\frac{1}{2}re^{\lambda}(\epsilon+p)H_{0}+re^{\lambda-\nu/2}\left(X+\frac{1}{2r^{2}}e^{\nu/2}T_{1}\right)+\left[\frac{1}{2}(\lambda^{\prime}-\nu^{\prime})-\frac{\ell+1}{r}\right]T_{2}+\left[\frac{2n_{\ell}e^{\lambda}\mu}{r}-e^{\lambda-\nu}r\omega^{2}(\epsilon+p)\right]V+e^{\lambda/2}p^{\prime}W\,.$ (76)
@@
$\left[\frac{re^{-\lambda}}{2}\left(r\nu^{\prime}-2\right)+(n_{\ell}+1)r\right]H_{0}$ $=r^{2}e^{-\lambda}\left[\omega^{2}re^{-\nu}-\frac{n_{\ell}+1}{2}\nu^{\prime}\right]H_{1}$
@@
$\frac{1}{3}e^{-\nu/2}\mu r^{2}X+\frac{1}{4}\Gamma_{1}pT_{1}$ $=\frac{\mu\Gamma_{1}p}{2}\left[2e^{-\lambda/2}W-r^{2}K+2(n_{\ell}+1)V\right].$ (79)
@@ A.2 Joint and boundary conditions
The solutions must remain regular at the center of the star, necessitating a Taylor expansion of the perturbation variables near the origin to approximate the ODE solutions. It has been shown in Ref. [37] that not all the quantities are independent. Once $K(r=0)$ and $W(r=0)$ are specified, all other coefficients in the Taylor expansion can be uniquely determined (see details in Ref. [89, pp. 27–30, 59–60]). The surface of the star is a matter-vacuum interface. If a two-layer NS with a solid surface is considered, the boundary conditions at the surface are